32Men working at the rate of 9hrs aday can complete apiece ofwork in 7 days.how many more men working at therate of 8hrs a day would complete the work in 6days

1
Expert's answer
2022-03-08T05:53:02-0500

Number of men required to complete the work in 6 days "=\frac{9}{8}\times \frac{7}{6} \times 32=42"

Extra men required "=42-32=10men"